Let G be a semisimple linear algebraic group over a field k and let X be a projective homogeneous G-variety. The variety X is called generically cellular if G splits over the field of rational functions k(X). For example, a Pfister quadric is a generically cellular variety.

In the talk I intend to classify such varieties and give applications to the Rost invariant.